• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name N Toro

    Log entry time 00:01:30 on July 7,2010

    Entry number 330549

    Followups:

    keyword=optics runs / last hr of swing shift

    Shift summary for remainder of swing shift, and sieve-in runs for owl shift
    (Kalyan had to leave, Natalia taking over at 23:18)

    ===Sieve-In Optics Runs===
    23:18 -- back to beam permit.
    During access:
    * Sieve inserted on L and R
    * Bogdan modified T6 temporarily to HRS-R S2m+ HRS-R GC (positron singles trigger)

    23:21 -- Run 1810 = R-arm sieve-in optics. 20uA on Ta, raster off. 1M events on modified T6 trigger (could not end run, EB2 was not responding -- may be corrupted)

    23:35 We disconnected and reconnected coda, and took a short test run (1811, 50K events) which seemed to work.

    23:36 -- Runs 1812, 1813 = same settings as 1810, 1M events

    00:00 -- Runs 1814, 1815 = L-arm sieve-in optics. 3uA on Ta, raster off. 1M events per run on T5 trigger